Q(3)It is known that only 60% of a defected computer can be   

   repaired. A sample of 15 computers is selected randomly,   

   find the probability that;

 (i) None can be repaired  ​        

 (ii) 5 or less can be repaired ​     

(iii) At least 3 computers can be repaired. 

 



1
Expert's answer
2021-08-10T11:51:35-0400

Laptop, which can be repaired "=60\%"

Total number of computer = 15

Total number of laptop which can be repaired "= 60\times \frac{15}{100} =9"

Total number of laptop which can not be repaired "= 6"

a) Probability, to find the laptop which can not be prepared "=\frac{6}{15}=\frac{2}{5}"

b) The required probability "=\frac{C^9_5}{C^{16}_5}=\frac{9\times 8\times 7\times 6}{16\times15\times14\times 13}"

"=\frac{9}{130}"

c) The required probability "=\frac{C^9_3}{C^{16}_3}=\frac{9\times 8\times 7}{16\times 15\times 14}"

"=\frac{3}{20}"